# Enhanced pedestal transport driven by edge collisionality on Alcator C-Mod and its role in regulating H-mode pedestal gradients

### v1

@@ 5 Neutral modeling and interpretive transport solutions with SOLPS-ITER
The technique outlined in Section 2 is powerful in that it can be employed across an entire database of shots with high quality Lyα and ETS data. They depend on 1D measurements, however, and any conclusions about transport hinge on assumptions about the poloidal distribution of neutrals. If $\Gamma_{D}$ estimated at the outer midplane (OMP) ($\Gamma_{D}^{\mathrm{OMP}}$) is significantly different from the flux-surface averaged $\Gamma_{D}$ ($\Gamma_{D}^{\mathrm{FSA}}$) the experimentally inferred $D_{\mathrm{eff}}$ may not be wholly indicative of transport in the pedestal across its poloidal extent. Estimates of $D_{\mathrm{eff}}^{\mathrm{ped}}$ assume that variation in the poloidal fueling profiles between shots is small and that generally, this peaks at or near the poloidal location observed by the Lyα camera (here, the OMP). Furthermore, without some interpretive power balance analysis using a transport code, we can only investigate the particle transport and know little about changes to the electron or ion thermal transport. To gain insight into thermal transport, validate our inferences of particle transport, and learn about poloidal distribution of neutrals, we choose certain discharges from these experiments to simulate with SOLPS-ITER, a code suite used extensively for edge modeling of plasma and neutrals. SOLPS-ITER couples a 2D multi-fluid plasma transport code, B2.5, with a 3D kinetic Monte Carlo neutral transport solver, EIRENE [51]. In the current study, we focus primarily on the OMP and constrain the plasma and neutral models there. The simulations are performed for only deuterium, D, ion species.
@@ 5.1.1 Boundary conditions
For the plasma solution, we attempt to reproduce $n_{e}$ and $T_{e}$ from the top of the pedestal to the near-SOL using profile fits to the ETS measurements. This requires choosing appropriate boundary conditions (BC) and cross-field transport coefficients. At the core boundary, we prescribe input power and particle fluxes. For the thermal channel, we use our estimate for $P_{\mathrm{net}}$ and assume energy equipartition, providing $\frac{1}{2}P_{\mathrm{net}}$ to each the electron and the ion populations. In reality, $P_{\mathrm{net}}$ is an estimate of the net power crossing the separatrix, which in these simulations, is radially outside the core boundary. Applying it at the innermost flux-surface of the B2.5 plasma grid assumes negligible radiated power in the plasma annulus about 1 cm thick between this flux-surface and the separatrix. This could yield an underestimation of the actual power crossing that boundary. For the particle channel, we set the particle flux, $\Gamma=0$. Alcator C-Mod had no core particle source, neither via neutral beam injection (NBI), nor pellet-fueling. Any non-zero $\Gamma$ at the innermost boundary of the B2.5 grid comes entirely from neutrals recycled at its walls, assumed to be negligible. This is checked a posteriori to be a valid assumption, as shown below in Section 5.2.
Although these discharges all have the lower X-point active, they are all close to double-null (DN), with $\Delta R_{\mathrm{sep}}\approx 5$ mm. Accurately capturing the physics of the far-SOL required using a disconnected DN (DDN) grid, yielding four targets in the simulation domain. At each of these, sheath BCs are used, set via the usual modified Bohm criterion. The final boundaries left to define BCs for are the SOL and PFR boundaries. The BCs in both are set to the “leakage” type, instead of the more conventional “decay length” type. The choice for these BCs is important for reproducing the details of the neutral profiles, and are thus outlined in Section 5.2. Finally, flux limiters for the thermal parallel transport are included in the simulations, using a value of 0.15 for both ions and electrons.
@@ 5.1.2 Transport coefficients
For the plasma transport, we use a purely diffusive model across the three transport channels. For particles, we only assume diffusion, specifying a particle density diffusion coefficient, $D_{n}$. For energy, we use thermal diffusivity coefficients, $\chi_{e}$ for the electrons, and $\chi_{i}$ for the ions. To reproduce the steep gradients in these H-mode edges, it is necessary to apply a transport well by using a radially-varying profile across the three channels. We ignore poloidal variation of transport coefficients in the core and specify transport profiles that reproduce $n_{e}$ and $T_{e}$ measurements from ETS at the OMP. This is likely a source of error, as previous work on C-Mod has shown that ballooning-like transport drives larger plasma fluxes to the outboard rather than inboard side of the core [36]. We further assume negligible poloidal variation in radial transport along open field lines in the SOL, another potentially large simplification. In other words, the same radial SOL transport coefficient profile at the midplane is used at the divertor targets. Finally, we use flat transport coefficients in the private-flux region (PFR): $D_{n}$ = 5 $\times$ 10-2 m-2s-1 and $\chi_{e}=\chi_{i}$ = 7 $\times$ 10-1 m-2s-1.
Figure 11: Electron density (top) and particle diffusion coefficient (bottom) from a converged SOLPS simulation. The top plot shows the experimntal ETS data points and their errorbars, as well as the best mtanh fit (dashed blue line) computed from a least-squares minimization procedure. The orange line shows the resulting simulated $n_{e}$ after a set of iterations to determine the $D(r)$ profile, plotted below that would best reproduce the experimental $n_{e}(r)$.
Figure 12: The neutral density profile used to constrain the EIRENE solution, in log scale. In blue are the points, along with the error bars, corresponding to the measured $n_{e}$ and $T_{e}$ from ETS, onto which $\epsilon_{\mathrm{Ly}_{\alpha}}$ is interpolated to experimentally infer $n_{0}$. In orange is the simulated profile computed by EIRENE and interpolated onto the B2.5 plasma grid. The orange profile has been tuned using the procedure in Section 5.2 to get as best a visual match in the pedestal region as possible.
@@
1070821003 1.4 240 2.4 3.7 40
1070821004 2.0 280 2.5 3.0 28
1070821009 2.9 480 2.2 2.6 23
1070821008 3.4 700 2.3 2.5 21
Table 2: Relevant parameters pertaining to the four selected discharges to be simulated using SOLPS-ITER. As before, the discharges at low $P_{net}$ have high $n_{e}$, low $T_{e}$, and high $n_{0}$. The opposite is true for the discharges at high $P_{\mathrm{net}}$.
Figure 13: Pedestal plasma solutions (top) and transport coefficient profiles (bottom) for the four simulated discharges, tabulated in Table 2. Quantities are shown for the particle transport channel (left) and for the electron thermal transport channel (right). $n_{e}$ and $D$ profiles are largely self-similar, with differences in magnitude similar to those gauged experimentally. $T_{e}$ and $\chi_{e}$ profiles are not, varying significantly in height and in width. In the absence of $T_{i}$ measurements, $\chi_{i}$ is used simply to enforce $T_{e}=T_{i}$ and thus, physical significance is uncertain.
To determine transport profiles for the plasma in the MC, we employ an iterative approach. This technique is based on a diffusive transport ansatz, and has been used to reproduce TS profiles on NSTX and DIII-D [52, 53]. First, the code is run to convergence using arbitrary $\{D_{n},\chi_{e},\chi_{i}\}$ profiles. We evaluate the particle flux, electron heat flux, and ion heat flux from the converged solution. We then use experimentally measured gradients from ETS to calculate a new set of transport coefficients. These new transport coefficients modify simulated gradients, a function of the simulated fluxes, in order to match experiment. This calculation is done at every radial cell, yielding three radially-varying profiles, i.e $\{D_{n}(r),\chi_{e}(r),\chi_{i}(r)\}$. These new coefficients calculated at iteration $j+1$ are given in Equations 8 and 9 and depend on the fluxes at the previous iteration $j$.
@@
In Equations 8 and 9, $q_{e,i}$ is the heat flux either for the electrons or ions. Equation 9 is actually two equations, one for electrons and one for ions. Since ion temperature, $T_{i}$, measurements are unavailable for these cases we assume $T_{e}=T_{i}$, and use $\nabla T_{e}^{exp}$ in lieu of $\nabla T_{i}^{exp}$ in the calculation of $\chi_{i}$. Earlier work comparing $T_{i}$ to $T_{e}$ showed that the temperature ratio, $\tau_{i}=T_{i}/T_{e}$, was not far from one at the pedestal, but could be as high as six at the separatrix and in the SOL [54]. $\tau_{i}$ was found to be most clearly dependent on divertor collisionality - the more collisional the divertor, the more equilibrated $T_{i}$ and $T_{e}$. Given that the EDA plasmas in the current study were very collisional, $\tau_{i}=1$, while likely an under-estimation, may not be wholly inaccurate, at least through the pedestal. Recent simulations of DIII-D have tested various approaches to setting the $T_{i}$ temperature. They show sensitivity of outputs, namely increased neutral pressure in the PFR at higher values of $T_{i}$ upstream. This motivates more detailed study of the sensitivity of upstream $n_{0}(r)$, specifically in the pedestal, to choice of $T_{i}$ [53].
Figure 11 shows an example of experimental ETS data and the corresponding mtanh fit. Plotted also is the simulated $n_{e}$ profile as well as the corresponding $D_{n}$ profile used to match the experimental $\nabla n_{e}$ from the mtanh fit. Using this iterative transport solving scheme, between $25-35$% of input power is lost radially across the outer grid boundary. The iterative scheme is a computationally expensive process, so we run the code without including fluid drifts or currents.
@@ 5.2 Neutrals constraints
A consequence of Equation 8 is that without a constraint for $\Gamma^{\mathrm{SOLPS}}$, any number of values of $D_{n}$ can reproduce the requisite $\nabla n^{\mathrm{exp}}$. In other words, $n_{e}$ on its own is insufficient to constrain $\Gamma^{\mathrm{SOLPS}}$ and thus, to find the correct $D_{n}$. As shown in Section 2, knowing $n_{0}$ or equivalently $S_{\mathrm{ion}}$ fixes $\Gamma_{D}$, which can constrain the EIRENE calculation. Therefore, only if we can simultaneously match $n_{e}$, $T_{e}$, and $n_{0}$ at the OMP can we confidently make a conclusion about $D_{n}$, as well as $\chi_{e}$ [27].
@@
Recent work on Alcator C-Mod has helped validate the $n_{0}$ calculation in EIRENE using the same LYMID camera [27]. A match to experimental $n_{0}$ was achieved through minimal tuning and across confinement modes, using leakage BCs in the SOL and PFR, as mentioned in Section 5.1.1. Rather than specifying a decay length given by $n_{e}$ and $D_{n}$ at the outermost grid cell, leakage BCs insist that the particle flux returning from the vessel walls be proportional to the local sound speed ($c_{s}$) and $n$, i.e. $\Gamma\sim\sqrt{\frac{T_{e}+T_{i}}{m_{i}}}n$. The current work finds that the proportionality constant for the particle channel, the particle leakage coefficient ($\alpha_{n}$), sets the magnitude of $\Gamma$ (and by consequence, $n_{0}$). This BC is thus used to vertically shift $n_{0}$, attempting primarily to match the magnitude of $n_{0}$ in the pedestal. While $\alpha_{n}$ changes the magnitude of $n_{0}$ it keeps the neutral gradient scale length ($L_{n_{0}}$) in the pedestal ($L_{n_{0}}^{\mathrm{ped}}$) relatively fixed. Previous studies of neutral penetration shows that $L_{n_{0}}$ depends strongly on $n_{e}$ [15]. Since we are seeking a particular $n_{e}$, it is not a free parameter, and we instead mobilize the thermal leakage coefficients for electrons ($\alpha_{T_{e}}$) and ions ($\alpha_{T_{i}}$) to tune $L_{n_{0}}^{\mathrm{ped}}$. These essentially modify $q_{e,i}$, which indirectly affects $\Gamma_{D}$, essentially modifying the slope of $n_{0}$. The coefficients $\alpha_{n}$, $\alpha_{T_{e}}$, $\alpha_{T_{i}}$} thus become the toolkit needed to match both the magnitude and slope of $n_{0}$ across the pedestal, given a particular B2.5 solution. Updating the EIRENE solution also affects B2.5 by modifying the sources for the plasma transport equations. This essentially adds an outer loop to the iterative process described in Section 5.1. Figure 12 shows an example of the $n_{0}$ profile, calculated by EIRENE and interpolated onto the B2.5 grid, which results from this tuning process and which corresponds to the case shown in Figure 11.
@@
For these discharges, the simulation domain is treated as an isolated system, with no injection or removal of particles. Since Alcator C-Mod had molybdenum walls which did not readily absorb deuterium, we set the albedo (R) to unity at all vessel walls, including targets. As mentioned in Section 5.1 we impose $\Gamma=0$ at the core boundary. Recalling that these H-modes were not actively puffed (see Section 2), we also do not inject gas into the simulation domain via gas puffing. Particle balance is therefore achieved in these simulations simply by manipulating the initial particle inventory.
@@ 5.3 Transition in particle and thermal transport
Given the computational cost of these simulations, we select four discharges from the power scan to model, covering a range in $P_{\mathrm{net}}$ and $\nu^{*}$. Table 2 shows the four discharges to be simulated. Two of these are below the critical $P_{\mathrm{net}}$ and two are above. Key parameters of the simulated discharges are listed in Table 2. Figure 13 shows simulated $n_{e},T_{e},D_{n}$ and $\chi_{e}$ profiles across the pedestal and into the SOL for the four simulated discharges. The $n_{e}$ and $T_{e}$ profiles shown here all closely match the fitted experimental profiles, within fit error bars. As expected, the high $P_{\mathrm{net}}$ discharges have prominent $T_{e}$ pedestals and slightly lower $n_{e}$ pedestals. As $P_{\mathrm{net}}$ drops, $T_{e}^{\mathrm{ped}}$ plummets and $n_{e}^{\mathrm{ped}}$ grows to the saturated level seen earlier, in Figure 5. At this scale, it is hard to discern variation in separatrix conditions. At the pedestal top, however, it is evident that the green curves are at considerable higher $\nu^{*}$ than the purple curves, especially given the $1/T_{e}^{2}$ dependence in $\nu^{*}$.
@@
The bottom panels of Figure 13 show the transport profiles found to reproduce $n_{e},T_{e}$ and $n_{0}$. In the pedestal, $D_{n}(r)$ shifts down and slightly radially inwards as $P_{\mathrm{net}}$ increases. In other words, dropping power shifts the characteristic H-mode transport well out, closer to open field lines, and to larger values of $D$. There do not, however, appear to be drastic changes to the structure of the profile within the pedestal. This is not surprising given the apparent self-similarity of the $n_{e}$ profiles and the small variation in profile gradients. $\chi_{e}$, on the other hand, clearly changes both in magnitude and in profile structure across the $P_{\mathrm{net}}$ scan. The rise in transport observed in the particle channel is even larger in the electron thermal channel. The purple curves have much higher $\chi_{e}$ across the pedestal than the green curves. They also appear to shift slightly radially inward, in the opposite direction of $D_{n}$ at low $P_{\mathrm{net}}$. It follows then that the relative shift between $n_{e}$ and $T_{e}$ shown in Section 4 can be explained by the spatial decoupling of the particle and electron thermal transport shear layers in low $P_{\mathrm{net}}$ plasmas. More notably, however, is the significant widening of $\chi_{e}(r)$ as $P_{\mathrm{net}}$ drops. The $T_{e}$ pedestals at low $P_{\mathrm{net}}$, shown in the top right panel, are considerably wider than those at high $P_{net}$. At the separatrix, $D\sim 6\times 10^{-2}$ m2s-1 and $\chi_{e}\sim 3\times 10^{-1}$ m2s-1 for all values of $P_{\mathrm{net}}$.
@@
In Figure 14, we plot the simulation results onto the experimentally inferred $D_{\mathrm{eff}}$, as a function of $\nu^{*}_{\mathrm{sep}}$. Indeed the simulated $D_{\mathrm{eff}}$ from SOLPS fall within the scatter of the experimental data, trending upwards with collisionality. This lends confidence to our ability to experimentally infer $D_{\mathrm{eff}}$ from a 1D neutral measurement localized at the OMP on C-Mod. Of course, as different machines may have different poloidal neutral distributions, this is not in general the case across machines. Figure 14 shows that it is actually $\chi_{e}$ that undergoes the most dramatic transition at a critical collisionality, $\nu^{*}_{\mathrm{sep,crit}}=25$. Interestingly enough, at $\nu^{*}_{\mathrm{sep}}<\nu^{*}_{\mathrm{sep,crit}}$, $D_{n}\sim\chi_{e}$. At $\nu^{*}_{\mathrm{sep}}>\nu^{*}_{\mathrm{sep,crit}}$, however, $D_{n}<<\chi_{e}$. Recent computational work using pedestal gyrokinetics has indicated that the ratio of diffusivities in the pedestal in different transport channels may be indicative of different transport drives [55]. Indeed this may be consistent with the earlier assertion of transition in the character of turbulence from Figure 10.
@@
Figure 14: Transport coefficients, inferred experimentally (diamonds) or computationally (stars) against separatrix collisionality. The blue diamonds are the experimentally inferred particle transport coefficients. The yellow stars are the same coefficients, but determined computationally using Equation 8 iteratively. The red stars are the electron thermal transport coefficients determined via the simulation, using the electron component of Equation 9. $D_{\mathrm{eff}}$ coefficients are extracted at mid-$n_{e}$ pedestal, while $\chi_{e}$ is extracted at mid-$T_{e}$ pedestal. $D_{\mathrm{eff}}^{\mathrm{SOLPS}}$ fall within the scatter of $D_{\mathrm{eff}}^{\mathrm{exp}}$, while $\chi_{e}^{\mathrm{SOLPS}}$ is larger at high $\nu^{*}_{\mathrm{sep}}$ but close in magnitude at low $\nu^{*}_{\mathrm{sep}}$

### v2

@@ 5 Neutral modeling and interpretive transport solutions with SOLPS-ITER
The technique outlined in Section 2 is powerful in that it can be employed across an entire database of shots with high quality Lyα and ETS data. These are 1D measurements, however, so any conclusions about transport hinge on assumptions about the poloidal distribution of neutrals. If $\Gamma_{D}$ estimated at the outer midplane (OMP), $\Gamma_{D}^{\mathrm{OMP}}$, is significantly different its flux-surface averaged value, $\Gamma_{D}^{\mathrm{FSA}}$, the experimentally inferred $D_{\mathrm{eff}}$ may not be wholly indicative of transport in the pedestal across its poloidal extent. Experimental estimates of $D_{\mathrm{eff}}$ made here assume that variation in the poloidal fueling profiles between shots is small and that generally, this peaks at or near the poloidal location observed by the Lyα camera (here, the OMP). Furthermore, without some interpretive power balance analysis using a transport code, we can only investigate the particle transport and know little about changes to the electron or ion thermal transport. To gain insight into thermal transport, compare with our inferences of particle transport from 1D analysis, and learn about poloidal distribution of neutrals, we choose certain discharges from these experiments to simulate with SOLPS-ITER, a code suite used extensively for edge modeling of plasma and neutrals. SOLPS-ITER couples a 2D multi-fluid plasma transport code, B2.5, with a 3D kinetic Monte Carlo neutral transport solver, EIRENE [58]. In the current study, we focus primarily on the OMP and constrain the plasma and neutral models there. The simulations are performed for only deuterium, D, ion species.
@@ 5.1.1 Boundary conditions
For the plasma solution, we reproduce $n_{e}$ and $T_{e}$ from the top of the pedestal to the near-SOL using profile fits to the ETS measurements. At the core boundary, we prescribe input power and particle fluxes. We take $P_{\mathrm{net}}$ from experiment and assume energy equipartition, providing $\frac{1}{2}P_{\mathrm{net}}$ to each the electron and the ion populations. While the exact partition of energy in the edge can vary in different plasmas, the assumption of equal heat fluxes falls within a range of values estimated from TRANSP runs of a number of C-Mod plasmas [59]. Applying $P_{\mathrm{net}}$ at the innermost flux surface of the B2.5 plasma grid rather than at the separatrix assumes negligible radiated power in the plasma annulus about 1 cm thick between this flux surface and the separatrix. For the particle flux, we set $\Gamma=0$. Alcator C-Mod had no core particle source, neither via neutral beam injection (NBI), nor pellet-fueling. Any non-zero $\Gamma$ at the innermost boundary of the B2.5 grid comes entirely from neutrals recycled at its walls, assumed to be negligible. This is checked a posteriori to be a valid assumption, as shown below in Section 5.2.
Because of the small $\Delta R_{\mathrm{sep}}\approx 5$ mm, a disconnected DN grid is used, yielding four targets in the simulation domain. At each of these, sheath BCs are used, set via the usual modified Bohm criterion. For the SOL and private flux region (PFR), BCs are set to the “leakage" type, instead of the more conventional “decay length" type. The choice for these BCs is important for reproducing the details of the neutral profiles and are outlined in Section 5.2. Finally, motivated by a study for the importance of using heat flux limiters to match simulations to experimental C-Mod divertor temperature measurements [60], flux limiters for the thermal parallel transport are included in the simulations, using a value of 0.15 for both ions and electrons.
@@ 5.1.2 Transport coefficients
For particle transport, we assume purely diffusive transport, specifying the particle density diffusion coefficient, $D_{n}$. For energy, we use thermal diffusivity coefficients, $\chi_{e}$ for the electrons, and $\chi_{i}$ for the ions. To reproduce the steep gradients in these H-mode edges, it is necessary to apply a transport well by using a radially-varying profile in all three channels. We ignore poloidal variation of transport coefficients in the core and specify transport profiles to reproduce $n_{e}$ and $T_{e}$ measurements from ETS, at the OMP. This is likely a source of error, as previous work on C-Mod has shown that ballooning-like transport drives larger plasma fluxes to the outboard rather than inboard side of the core [43]. As a result, the $D_{\mathrm{eff}}$ reported in this work may represent an underestimation of the transport on the LFS and an overestimation on the HFS. We further assume negligible poloidal variation in radial transport along open field lines in the SOL, such that the same transport coefficient profile at the midplane SOL is used at the divertor targets. Finally, we use flat transport coefficients in the PFR: $D_{n}$ = 5 $\times$ 10-2 m-2s-1 and $\chi_{e}=\chi_{i}$ = 7 $\times$ 10-1 m-2s-1.
Figure 10: Electron density (top) and particle diffusion coefficient (bottom) from a converged SOLPS simulation. The top plot shows the experimental ETS data points and their errorbars in open, gray circles, as well as the best mtanh fit (dashed black line) computed from the least-squares minimization procedure outlined in Section 2.3. The orange line shows the resulting $n_{e}$ after a set of iterations to determine the $D(\psi_{n})$ profile, plotted below, that would best reproduce the experimental $n_{e}(\psi_{n})$, while simultaneously matching $n_{0}(\psi_{n})$.
Figure 11: The neutral density profile used to constrain the EIRENE solution, in log scale. In open, gray circles are the points, along with the error bars, corresponding to the measured $n_{e}$ and $T_{e}$ from ETS, onto which $\epsilon_{\mathrm{Ly}_{\alpha}}$ is interpolated to experimentally infer $n_{0}$. The dashed black line represents the $n_{0}$ computed from the experimental profile fits. In turquoise is the simulated profile computed by EIRENE and interpolated onto the B2.5 plasma grid. This profile has been tuned using the procedure in Section 5.2 to get as best a visual match in the pedestal region as possible.
@@
1070821003 1.4 240 2.4 3.7 28
1070821004 2.0 280 2.5 3.0 24
1070821009 2.9 480 2.2 2.6 21
1070821008 3.4 700 2.3 2.5 13
Table 2: Relevant parameters pertaining to the four selected discharges to be simulated using SOLPS-ITER. As before, the discharges at low $P_{\mathrm{net}}$ have high $n_{e}$, low $T_{e}$, and high $n_{0}$. The opposite is true for the discharges at high $P_{\mathrm{net}}$.
Figure 12: Pedestal plasma solutions (top) and transport coefficient profiles (bottom) for the four simulated discharges tabulated in Table 2. Quantities are shown for the particle transport channel (left) and for the electron thermal transport channel (right). $n_{e}$ and $D$ profiles are largely similar in shape. $T_{e}$ and $\chi_{e}$ profiles are not, varying significantly in height and in width. In the absence of $T_{i}$ measurements, $\chi_{i}$ is used simply to enforce $T_{e}=T_{i}$ and is not shown, as physical significance is uncertain.
To determine transport profiles for the plasma in the main chamber and SOL, we employ an iterative approach used to reproduce TS profiles on NSTX and DIII-D [61, 62]. The approach uses a set of transport coefficient profiles at iteration $j$ to evaluate a new set of profiles at iteration $j+1$ by using the experimental gradients, according to the following,
@@
where $q_{e,i}$ is either the electron or ion heat flux. Equation 9 is solved for electrons and ions separately. Since ion temperature, $T_{i}$, measurements are unavailable for these cases, we assume $T_{e}=T_{i}$, and set $\nabla T_{i}^{\mathrm{exp}}=\nabla T_{e}^{\mathrm{exp}}$ in the calculation of $\chi_{i}$. Earlier work comparing $T_{i}$ to $T_{e}$ showed that the temperature ratio, $\tau_{i}=T_{i}/T_{e}$, was not far from unity at the pedestal, but could be as high as six at the separatrix and in the SOL [60]. $\tau_{i}$ was found to be most clearly dependent on divertor collisionality - the more collisional the divertor, the more equilibrated $T_{i}$ and $T_{e}$. Given that the EDA plasmas in the current study were very collisional, $\tau_{i}=1$, while likely an under-estimation, may not be wholly inaccurate, at least through the pedestal. Recent simulations of DIII-D have tested various approaches to setting the $T_{i}$ temperature. They show sensitivity of outputs, namely increased neutral pressure in the PFR at higher values of $T_{i}$ upstream. This motivates more detailed study of the sensitivity of upstream $n_{0}(\psi_{n})$, specifically in the pedestal, to choice of $T_{i}$ [62].
Figure 10 shows an example of experimental ETS data and the corresponding mtanh fit. Plotted also is the simulated $n_{e}$ profile as well as the corresponding $D_{n}$ profile used to match the experimental $\nabla n_{e}$ from the mtanh fit. Using this iterative transport solving scheme, between $25-35$% of input power exits the grid radially across the outer grid boundary, presumably hitting the first wall and not arriving at the divertor. Furthermore, for simplicity, the code is run without including fluid drifts or currents.
@@ 5.2 Neutrals constraints
A consequence of Equation 8 is that without a constraint for $\Gamma^{\mathrm{SOLPS}}$, any number of values of $D_{n}$ can reproduce the requisite $\nabla n^{\mathrm{exp}}$. In other words, $n_{e}$ on its own is insufficient to constrain $\Gamma^{\mathrm{SOLPS}}$ and thus, $D_{n}$. As shown in Section 2, knowing $n_{0}$ or equivalently $S_{\mathrm{ion}}$ fixes $\Gamma_{D}$, which does constrain the EIRENE calculation. Therefore, only if we can simultaneously match $n_{e}$, $T_{e}$, and $n_{0}$ at the OMP can we confidently make a conclusion about $D_{n}$, as well as $\chi_{e}$ [27], at least in a poloidally-averaged sense.
@@
Recent work on Alcator C-Mod has helped validate the $n_{0}$ calculation in EIRENE using the same LYMID camera [27] across confinement modes. As mentioned in Section 5.1.1, rather than specifying a decay length given by $n_{e}$ and $D_{n}$ at the outermost grid cell, we use leakage BCs. These insist that the particle flux returning from the vessel walls be proportional to the local sound speed, $c_{s}$, and $n$, i.e. $\Gamma\sim\sqrt{\frac{T_{e}+T_{i}}{m_{i}}}n$. The current work finds that the proportionality constant for the particle channel, the particle leakage coefficient, $\alpha_{n}$, sets the magnitude of $\Gamma$ (and by consequence, $n_{0}$). This BC is thus used to shift $n_{0}$ up and down in magnitude, attempting primarily to match the magnitude of $n_{0}$ in the pedestal. While $\alpha_{n}$ changes the magnitude of $n_{0}$ it keeps the neutral gradient scale length, $L_{n_{0}}$, in the pedestal, $L_{n_{0}}^{\mathrm{ped}}$, i.e. the shape of the neutral density profile, relatively fixed. Previous studies of neutral penetration shows that $L_{n_{0}}$ depends strongly on $n_{e}$ [15]. Since we are seeking a match to $n_{e}$, we must also match $L_{n_{0}}$, and we instead mobilize the thermal leakage coefficients for electrons, $\alpha_{T_{e}}$, and ions, $\alpha_{T_{i}}$, to tune $L_{n_{0}}^{\mathrm{ped}}$. These essentially modify $q_{e,i}$, which indirectly affects $\Gamma_{D}$ by modifying the shape of $n_{0}$. The coefficients $\alpha_{n}$, $\alpha_{T_{e}}$, $\alpha_{T_{i}}$} thus become the toolkit needed to match both the magnitude and slope of $n_{0}$ across the pedestal, given a particular B2.5 solution. Updating the EIRENE solution, however, also affects B2.5 by modifying the sources for the plasma transport equations. This essentially adds an outer loop to the iterative process described in Section 5.1.2. Figure 11 shows an example of the $n_{0}$ profile, calculated by EIRENE and interpolated onto the B2.5 grid, which results from this tuning process and which corresponds to the case shown in Figure 10.
@@
For these discharges, the simulation domain is treated as an isolated system, with no injection or removal of particles. Since Alcator C-Mod had molybdenum walls which did not readily absorb deuterium, we set the albedo to unity at all vessel walls, including targets. As mentioned in Section 5.1.1 we impose $\Gamma=0$ at the core boundary. Recalling that these H-modes were not actively puffed (see Section 2), we also do not inject gas into the simulation domain via gas puffing. Particle balance is therefore achieved in these simulations simply by manipulating the initial particle inventory.
@@ 5.3 Transition in particle and thermal transport
Given the computational cost of these simulations, we select four discharges from the power scan to model, covering a range in $P_{\mathrm{net}}$ and $\nu^{*}$. Two of these are below the critical $P_{\mathrm{net}}$ and two are above. Key parameters of the four simulated discharges are listed in Table 2. Figure 12 shows simulated $n_{e},T_{e},D_{n}$ and $\chi_{e}$ profiles across the pedestal and into the SOL for the four simulated discharges. The $n_{e}$ and $T_{e}$ profiles shown here all closely match the fitted experimental profiles, within fit error bars. As expected, the high $P_{\mathrm{net}}$ discharges have prominent $T_{e}$ pedestals and slightly lower $n_{e}$ pedestals. As $P_{\mathrm{net}}$ drops, $T_{e}^{\mathrm{ped}}$ plummets and $n_{e}^{\mathrm{ped}}$ grows to the saturated level seen earlier, in Figure 4. At this scale, it is hard to discern variation in separatrix conditions. At the pedestal top, however, it is evident that the green curves are at considerably lower $\nu^{*}$ than the purple curves, especially given the $T_{e}^{-2}$ dependence in $\nu^{*}$.
@@
The bottom panels of Figure 12 show the transport profiles found to reproduce $n_{e},T_{e}$ and $n_{0}$. In the pedestal, $D_{n}(\psi_{n})$ shifts down and slightly radially inwards only at the highest value of $P_{\mathrm{net}}$. There are no drastic changes to the structure of the profile in the steep-gradient region of the pedestal. This is not surprising given the apparent similarity in shape of the $n_{e}$ profiles and the small variation in profile gradients. $\chi_{e}$, on the other hand, clearly changes both in magnitude and in profile structure across the $P_{\mathrm{net}}$ scan. Indeed, the rise in transport observed in the particle channel is even larger in the electron thermal channel. The purple curves have much higher $\chi_{e}$ across the steep gradient region than the green curves. They also appear to shift slightly radially inward, in the opposite direction of $D_{n}$ at low $P_{\mathrm{net}}$. It follows then that the relative shift between $n_{e}$ and $T_{e}$ and weakening of $\alpha_{\mathrm{MHD}}$ mentioned in Section 4 could also be explained by the spatial decoupling of the particle and electron thermal transport shear layers in low $P_{\mathrm{net}}$ plasmas. More notable, however, is the significant widening of $\chi_{e}(\psi_{n})$ as $P_{\mathrm{net}}$ drops. The $T_{e}$ pedestals at low $P_{\mathrm{net}}$, shown in the top right panel, are considerably wider than those at high $P_{\mathrm{net}}$. At the separatrix, $D\sim 6\times 10^{-2}$ m2s-1 and $\chi_{e}\sim 3\times 10^{-1}$ m2s-1 for all values of $P_{\mathrm{net}}$. This change in $\chi_{e}$ best explains the weakening of the $\nabla T_{e}$ observed in Section 4.
@@
In Figure 13, we plot the simulated $D_{n}$ and $\chi_{e}$ onto the experimentally inferred $D_{\mathrm{eff}}$, as a function of $\nu^{*}_{\mathrm{sep}}$. Indeed the simulated $D_{\mathrm{eff}}$ from SOLPS fall within the scatter of the experimental data, trending upwards with collisionality. This acts as a confirmation that the use of 1D analysis of plasma and neutral profiles as described in Section 2 yields a value for $D_{\mathrm{eff}}$ consistent with the poloidally-averaged one used in the SOLPS simulations. Of course, as different machines may have different poloidal neutral distributions, this confirmation is not in general the case across machines. Figure 13 shows that it is actually $\chi_{e}$ that undergoes the most dramatic transition at a critical collisionality, $\nu^{*}_{\mathrm{sep,crit}}\approx 22$. Interestingly enough, at $\nu^{*}_{\mathrm{sep}}<\nu^{*}_{\mathrm{sep,crit}}$, $D_{n}\sim\chi_{e}$. At $\nu^{*}_{\mathrm{sep}}>\nu^{*}_{\mathrm{sep,crit}}$, however, $D_{n}<<\chi_{e}$. Recent computational work using pedestal gyrokinetics has indicated that the ratio of diffusivities in the pedestal in different transport channels may be indicative of different transport drives [63]. Indeed this may be consistent with the earlier assertion of transition in the character of turbulence from Figure 9.
@@
Figure 13: Transport coefficients, inferred experimentally (diamonds) or computationally (stars) against separatrix collisionality. The blue diamonds are the experimentally inferred particle transport coefficients. The yellow stars are the same coefficients, but determined computationally using Equation 8 iteratively. The red stars are the electron thermal transport coefficients determined via the simulation, using the electron component of Equation 9. $D_{\mathrm{eff}}$ coefficients are extracted at mid-$n_{e}$ pedestal, while $\chi_{e}$ is extracted at mid-$T_{e}$ pedestal. $D_{\mathrm{eff}}^{\mathrm{SOLPS}}$ fall within the scatter of $D_{\mathrm{eff}}^{\mathrm{exp}}$, while $\chi_{e}^{\mathrm{SOLPS}}$ is larger at high $\nu^{*}_{\mathrm{sep}}$ but closer in magnitude at low $\nu^{*}_{\mathrm{sep}}$.